# Kilonova Light Curve Parameter Estimation Using Likelihood-Free Inference

### v2

@@ Acknowledgements
We thank the NMMA team for their invaluable discussions and suggestions for additional testing of the model framework and the incorporation of the model into the NMMA codebase. We thank Peter Pang for reviewing the code. We thank Ethan Marx, Will Benoit, and Alec Gunny for their helpful comments. This work is supported by NSF HDR Institute Grant PHY-2117997, “Accelerated AI Algorithms for Data-Driven Discovery.”
@@
The data processing code, neural network and normalizing flow code, training code, and model weights are available at: https://github.com/nuclear-multimessenger-astronomy/nmma. Through this package, command line inference using LFI is available. The training data is also available for use at: https://doi.org/10.5281/zenodo.11640528.
@@ Appendix A Hyperparameter Tuning
Tables 3 and 4 display the ten best hyperparameter tuned runs for the embedding network and the normalizing flow. Trial 2 and Trial 40 are selected as the respective best trials based on the lowest validation loss achieved at the termination of the run. These runs are shown in figure 8, which displays the training loss curves for the ten best trials for the embedding and for the normalizing flow. To increase the efficiency of hyperparameter searches, we leverage the Async HyperBand Scheduler (ASHA) from Ray. This scheduling algorithm uses the user-selected optimization metric (here, minimizing the validation loss) to identify poorly performing runs and terminate them early (Li et al., 2018). In this way, we can efficiently search over the parameter spaces in Table 2 for our models.
@@
Appendix B Model Resilience
As we intend for our framework to be used with a variety of optical instruments, it is imperative that the model can adjust to several cadences. For our first test, we discard every other detection and interpolate new data points in their stead. Again, we find that the model has sufficient information to generate an inference with high accuracy. Figure 9 displays a sample posterior in yellow alongside its respective light curve with the interpolated values. We observe alignment between the two.
Figure 9: Left: Original light curve data plotted in color with the interpolated points plotted as black stars for all three ZTF bands. Right: Posteriors for a light curve with $\log_{10}(M_{ej}/M_{\odot})$$=-1.1045$, $\log_{10}(V_{ej}/c)$$=-1.0095$, and $\log_{10}(X_{lan})$$=-6.2940$. Blue represents the flow results on the original data, and yellow with every other data point interpolated. The two posteriors overlap, indicating that interpolating these missing values has little effect on the resulting distribution.
This test indicates that our model has enough flexibility to accommodate different observational cadences to perform accurate parameter estimation.

### v3

@@ Acknowledgements
We thank the NMMA team for their invaluable discussions and suggestions for additional testing of the model framework and the incorporation of the model into the NMMA codebase. We thank Peter Pang for reviewing the code. We thank Ethan Marx, Will Benoit, and Alec Gunny for their helpful comments. This work is supported by NSF HDR Institute Grant PHY-2117997, “Accelerated AI Algorithms for Data-Driven Discovery.” M.W.C also acknowledges support from the National Science Foundation with grant numbers PHY-2308862 and PHY-2409481. We thank Michelle Ntampaka of the Space Telescope Science Institute for insightful discussions into building robust neural networks.
@@
The data processing code, neural network and normalizing flow code, training code, and model weights have been integrated into the NMMA codebase at: https://github.com/nuclear-multimessenger-astronomy/nmma. Through this package, command line inference using LFI is available. Previous and ongoing code development is located at: https://github.com/malinadesai/Kilo. The training data is also available for use at: https://doi.org/10.5281/zenodo.11640528.
@@ Appendix A Hyperparameter Tuning
Tables 5 and 6 display the ten best hyperparameter tuned runs for the embedding network and the normalizing flow. Trial 2 and Trial 40 are selected as the respective best trials based on the lowest validation loss achieved at the termination of the run. These runs are shown in Figure 8, which displays the training loss curves for the ten best trials for the embedding and for the normalizing flow. To increase the efficiency of hyperparameter searches, we leverage the Async HyperBand Scheduler (ASHA) from Ray. This scheduling algorithm uses the user-selected optimization metric (here, minimizing the validation loss) to identify poorly performing runs and terminate them early (Li et al., 2018). In this way, we can efficiently search over the parameter spaces in Table 3 for our models.
@@
Appendix B Further Resilience Testing
We investigated the results of interpolation to generate missing sections of data by removing four consecutive data points from each band near the peak and performing a linear interpolation over a 36-hour time span. The posteriors for the original and the interpolated light curves are well aligned. A sample case is shown in Figure 9 with the interpolated light curve shown on the left and its posterior shown in grey on the right. The original light curve posterior is shown in red. The two posteriors are aligned, providing a good indication that linear interpolation can be sufficient when the data has missing values. We also highlight that the true values of the light curve are correctly predicted by the model. Further work into interpolation schemes will be a useful avenue to increasing the accuracy of predicted points.
Figure 9: Left: Original light curve data plotted in color with the interpolated points plotted on top as black stars for all three ZTF bands. Right: Posteriors for the light curve pictured, with injection parameters $\log_{10}(M_{\rm{ej}}/M_{\odot})$$=-1.44$, $\log_{10}(V_{\rm{ej}}/c)$$=-0.81$, and $\log_{10}(X_{\rm{lan}})$$=-8.76$. Red represents the flow results on the original shifted light curve data, and grey represents the flow results for light curve data with four consecutive data points interpolated. The two posteriors mostly overlap, indicating that interpolating missing values has a minor effect on the resulting distribution.